Grade 12Algebra

Find
3.25 ÷ 0.5
Find
3.25 ÷ 0.5
Find
3.25 ÷ 0.5
Find
3.25 ÷ 0.5

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

We have,
= (325/100) ÷ (5/10)
= (325/100) × (10/5)
= (325 × 10)/ (100 × 5)
= (65 × 1)/ (10 × 1)
= 65/10
= 6.5

Thanks